OpenStudy (anonymous):

-4/√20 rationalize denominator

OpenStudy (jango_in_dtown):

\[\frac{ -4 }{ \sqrt{20} }=-4*\sqrt{20}/20=-\sqrt{20}/5\]

OpenStudy (jango_in_dtown):

just multiply the numerator and denominator by the same term for which it is in the sqrt

OpenStudy (anonymous):

in the what?

OpenStudy (dayakar):

\[\sqrt{20}=\sqrt{2*2*5}\]

OpenStudy (dayakar):

\[=2\sqrt{5}\]

OpenStudy (dayakar):

simplest form of denomnator

OpenStudy (dayakar):

now multiply numerator and denominator with sqrt 5 to rationalize the denominator

OpenStudy (dayakar):

\[\frac{ -4 }{ \sqrt{20} }= \frac{ -4 }{ 2\sqrt{5} }\]

OpenStudy (dayakar):

\[\frac{ -4 }{ 2\sqrt{5} }*\frac{ \sqrt{5} }{ \sqrt{5} }\]

OpenStudy (dayakar):

\[=\frac{ -4*\sqrt{5} }{ 2*5 }\]

OpenStudy (dayakar):

\[=\frac{ -2\sqrt{5} }{ 5 }\] cancellation with 2

OpenStudy (dayakar):

is it clear